# Predicting Video Slot Attention Queries from Random Slot-Feature Pairs

### v4

@@ Abstract
Unsupervised video Object-Centric Learning (OCL) is promising as it enables object-level scene representation and dynamics modeling as we humans do. Mainstream video OCL methods adopt a recurrent architecture: An aggregator aggregates current video frame into object features, termed slots, under some queries; A transitioner transits current slots to queries for the next frame. This is an effective architecture but all existing implementations both (i1) neglect to incorporate next frame features, the most informative source for query prediction, and (i2) fail to learn transition dynamics, the knowledge essential for query prediction. To address these issues, we propose Random Slot-Feature pair for learning Query prediction (RandSF.Q): (t1) We design a new transitioner to incorporate both slots and features, which provides more information for query prediction; (t2) We train the transitioner to predict queries from slot-feature pairs randomly sampled from available recurrences, which drives it to learn transition dynamics. Experiments on scene representation demonstrate that our method surpass existing video OCL methods significantly, e.g., up to 10 points on object discovery, setting new state-of-the-art. Such superiority also benefits downstream tasks like dynamics modeling. Our core source code and training logs are available as the supplement.
@@ Introduction
Unsupervised video object tracking. Unsupervised Video Object Tracking (VOT) resembles (li2019joint; yuan2020self; lai2020cvpr; wang2021unsupervised; xu2021iccv; shen2022cvpr; li2022cvpr) video OCL very much, because both are query-based and the query/slots should be maintained consistently across frames under self-supervision. However, VOT query only covers interested objects, while video OCL query covers all objects and the background. Essentially, unsupervised VOT only cares about temporal consistency of query representations while video OCL also demand slot representations to be informative enough to represent the scene. Video OCL slots, the sub-symbolic representations, can be directly employed in downstream scene representation and dynamics modeling, while VOT representations can not. Comparison with unsupervised VOT methods is of out scope.
@@
where $\boldsymbol{\phi}_{\mathrm{a}}$ is parameterized as typical Slot Attention (locatello2020slotattent); $s$ is the number of query vectors; $\boldsymbol{M}_{t}$ is binarized from $\boldsymbol{S}_{t}$’s attention maps along the slot dimension. Because Slot Attention is a kind of attention, $\boldsymbol{Q}_{t}$ and $\boldsymbol{S}_{t}$ always have identical tensor shapes. $\boldsymbol{M}_{t}$ can be used for object discovery, while $\boldsymbol{S}_{t}$ can be used for downstream scene representation and dynamics modeling.
@@
34.9
±0.3
DIAS 33.6
±0.4
49.3
±0.7
36.1
±1.4
35.2
±0.8
RandSF.Q 65.4 ±10.8 67.3 ±2.4 29.2 ±4.0 26.8 ±3.8 40.8
±3.3
77.7 ±1.3 26.8 ±1.0 25.2 ±1.0 48.1 ±1.0 59.9 ±1.9 42.4 ±0.2 41.6 ±0.2
Table 1: Object discovery on videos. Input resolution is $\times256$ ($\times224$); DINO2 ViT-S/14 is for encoding. On YTVIS, RandSF.Q surpasses SotA SlotContrast by 10+ points by ARI and ARIfg, and ˜7 points by mBO and mIoU.
@@
We comprehensively evaluate our method on object discovery, as well as downstream tasks of object recognition, visual question answering and object dynamics prediction. We also look into how well our method learns transition knowledge for query prediction. We conduct each experiment using three random seeds whenever it is applicable.
@@
Baselines. STEVE (singh2022steve) realizes the first Slot Attention-based OCL on videos. VideoSAUR (zadaianchuk2024videosaur) is the first method that utilizes vision foundation model and achieves competitive results in OCL on real-world complex videos. SlotContrast (manasyan2025slotcontrast) enables temporal consistency on long videos, setting new state-or-the-art (SotA) recently. DIAS (zhao2025dias), another state-of-the-art, supports OCL on both images and videos and we use the latter version without its slot pruning trick for fair comparison. Our method RandSF.Q is compared with these methods. Comparing with methods like SOLV (aydemir2023solv), which uses slot pruning, is unfair. So does comparing with SAVi (kipf2021savi) and SAVi++ (elsayed2022savipp), which are trained with external stronger supervision like optical flow and depth map.
Datasets. We evaluate these methods on both synthetic and real-world video datasets, following the experiments of SotA methods SlotContrast and DIAS. For synthetic video datasets, we use MOVi-C and MOVi-D of MOVi datasets series [^1] https://github.com/google-research/kubric/tree/main/challenge
@@
. These two subsets contain daily objects with incremental complex textures falling on complex background. These datasets are used for object discovery. For real-world video datasets, we use YouTube Video Instance Segmentation [^2] https://youtube-vos.org/dataset/vis (YTVIS) the high quality version [^3] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which consists of daily-life diverse and complex videos downloaded from YouTube. This dataset is used for both object discovery and object recognition, as well as object dynamics prediction. We also include the synthetic video-text dataset CLEVRER [^4] http://clevrer.csail.mit.edu, which consists of geometric objects and paired question texts, testing the multi-modal reasoning capability. This dataset is used for visual question answering.
@@
Object discovery. We compare our method RandSF.Q with baselines STEVE, VideoSAUR, SlotContrast and DIAS in terms of their object discovery performance on datasets MOVi-C, MOVi-D and YTVIS. This is basically an unsupervised object segmentation task. We use multiple recognized metrics for comprehensive measurement: Adjusted Rand Index (ARI) [^5] https://scikit-learn.org/stable/modules/generated/sklearn.metri
@@
As shown in Table 1, across all datasets and almost all metrics, our RandSF.Q defeats all the baselines. Specifically, on dataset YTVIS, RandSF.Q surpasses the latest SotA SlotContrast and DIAS by more than 10 points by metrics ARI and ARIfg, and up to 6 points by mBO and mIoU. In contrast, both SlotContrast and DIAS only boost their baseline VideoSAUR limitedly, less than 3 points by metrics mBO and mIoU, or even degrading by ARI and ARIfg.
Object recognition. We compare our method RandSF.Q with baseline SlotContrast in terms of their object recognition performance on dataset YTVIS. This consists of two sub-tasks, i.e., supervised classification and regression from slots. Some literature (locatello2020slotattent) names this as set prediction. We follow (seitzer2023dinosaur) to represent the dataset as slots and train a two-layer MLP to predict the object class and bounding box corresponding to each slot, supervised by the annotations of object class labels and bounding boxes in the dataset. We use top1 accuracy [^7] https://scikit-learn.org/stable/modules/generated/sklearn.metri
cs.accuracy_score.html
and R2 score [^8] https://scikit-learn.org/stable/modules/generated/sklearn.metri
cs.r2_score.html
as the class label classification and bounding box regression performance metrics.
@@
Dynamics Modeling
Better slot representation quality contributes to better dynamics modeling, which further benefits visual prediction, reasoning, planning and decision-making.
Object dynamics prediction. We focus on visual prediction as it has much less downstream factors that might confound with the results compared with other tasks. We compare our method RandSF.Q with baseline SlotContrast in term of their object dynamics prediction performance on dataset YTVIS. We follow the model architecturing settings from SlotFormer (wu2022slotformer). We first pre-train OCL models on YTVIS and freez them. Then we use the OCL model for representing video sample in YTVIS as slots, and input them into world model SlotFormer for predicting slots of future time steps. Like both (wu2022slotformer) and (manasyan2025slotcontrast), we do not measure the reconstructed pixel errors from predicted slots; instead, we measure the byproduct segmentation accuracy accumulated to current frame from predicted slots. The former may demonstrate good reconstruction performance but poor object dynamics. In contrast, the latter directly reflects how well each object being predicted and represented.
As shown in Figure 4, in the 15 rollout steps, our RandSF.Q always shows better object dynamics performance, measured by segmentation accuracy ARIfg and mBO, than baseline SlotContrast. Specifically, the significant performance gap between our method and the baseline is maintained throught the auto-regressive rollouting process of SlotFormer , the object-centric world model.
@@ Conclusion
Based on our two observations, we propose our method RandSF.Q. We realize informative query prediction by utilizing the next feature and effective query prediction learning by randomly sampling slot-feature pairs from available recurrences as transitioner inputs for query prediction. Our method show significant performance advantages over the latest SotA methods on scene representation, scene understanding and dynamics modeling. Our core claim is also verified through performance matrix analysis. Our work pushes the video OCL research greatly forward.

### v5

@@ Abstract
Unsupervised video Object-Centric Learning (OCL) is promising as it enables object-level scene representation and understanding as we humans do. Mainstream video OCL methods adopt a recurrent architecture: An aggregator aggregates current video frame into object features, termed slots, under some queries; A transitioner transits current slots to queries for the next frame. This is an effective architecture but all existing implementations both (i1) neglect to incorporate next frame features, the most informative source for query prediction, and (i2) fail to learn transition dynamics, the knowledge essential for query prediction. To address these issues, we propose Random Slot-Feature pair for learning Query prediction (RandSF.Q): (t1) We design a new transitioner to incorporate both slots and features, which provides more information for query prediction; (t2) We train the transitioner to predict queries from slot-feature pairs randomly sampled from available recurrences, which drives it to learn transition dynamics. Experiments on scene representation demonstrate that our method surpass existing video OCL methods significantly, e.g., up to 10 points on object discovery, setting new state-of-the-art. Such superiority also benefits downstream tasks like scene understanding.
Source Code, Model Checkpoints, Training Logshttps://github.com/Genera1Z/RandSF.Q
@@ Introduction
Unsupervised video object tracking. Unsupervised Video Object Tracking (VOT) resembles (li2019joint; yuan2020self; lai2020cvpr; wang2021unsupervised; xu2021iccv; shen2022cvpr; li2022cvpr) video OCL very much, because both are query-based and the query/slots should be maintained consistently across frames under self-supervision. However, VOT query only covers interested objects, while video OCL query covers all objects and the background. Essentially, unsupervised VOT only cares about temporal consistency of query representations while video OCL also demand slot representations to be informative enough to represent the scene. Video OCL slots, the sub-symbolic representations, can be directly employed in downstream scene representation and understanding, while VOT representations can not. Comparison with unsupervised VOT methods is of out scope.
@@
where $\boldsymbol{\phi}_{\mathrm{a}}$ is parameterized as typical Slot Attention (locatello2020slotattent); $s$ is the number of query vectors; $\boldsymbol{M}_{t}$ is binarized from $\boldsymbol{S}_{t}$’s attention maps along the slot dimension. Because Slot Attention is a kind of attention, $\boldsymbol{Q}_{t}$ and $\boldsymbol{S}_{t}$ always have identical tensor shapes. $\boldsymbol{M}_{t}$ can be used for object discovery, while $\boldsymbol{S}_{t}$ can be used for downstream scene representation and understanding.
@@
34.4
±0.2
RandSF.Qtsim 64.0 ±2.9 66.3 ±1.7 28.4 ±1.3 26.1 ±1.1 41.2
±2.2
72.0 ±1.1 27.1 ±0.9 25.4 ±0.9 46.8 ±0.7 60.7 ±1.7 41.5 ±0.2 40.6 ±0.1
RandSF.Qssc 65.4 ±10.7 67.4 ±2.1 29.2 ±3.8 26.8 ±3.7 41.6
±3.7
77.5 ±1.0 27.4 ±1.0 25.6 ±1.0 41.5 ±0.1 58.9 ±0.9 39.4 ±0.4 39.0 ±0.4
Table 1: Object discovery on videos. Input resolution is $\times256$ ($\times224$); DINO2 ViT-S/14 is for encoding. On YTVIS, RandSF.Q surpasses SotA SlotContrast by 10+ points by ARI and ARIfg, and ~7 points by mBO and mIoU. Note that tsim means using the time similarity loss from VideoSAUR (zadaianchuk2024videosaur), while ssc means using the slot-slot contrast loss from SlotContrast (manasyan2025slotcontrast).
@@
We comprehensively evaluate our method on object discovery, as well as downstream tasks of object recognition and visual question answering. We also look into how well our method learns transition knowledge for query prediction. We conduct each experiment using three random seeds whenever it is applicable.
@@
Baselines. STEVE (singh2022steve) realizes the first Slot Attention-based OCL on videos. VideoSAUR (zadaianchuk2024videosaur) is the first method that utilizes vision foundation model and achieves competitive results in OCL on real-world complex videos. SlotContrast (manasyan2025slotcontrast) enables temporal consistency on long videos, setting new state-or-the-art (SotA) recently. Our method RandSF.Q is compared with these methods. Comparing with methods like SOLV (aydemir2023solv), which uses slot pruning, is unfair. So does comparing with SAVi (kipf2021savi) and SAVi++ (elsayed2022savipp), which are trained with external stronger supervision like optical flow and depth map.
Datasets. We evaluate these methods on both synthetic and real-world video datasets, following the experiments of SotA method SlotContrast. For synthetic video datasets, we use MOVi-C and MOVi-D of MOVi datasets series [^1] https://github.com/google-research/kubric/tree/main/challenge
@@
. These two subsets contain daily objects with incremental complex textures falling on complex background. These datasets are used for object discovery. For real-world video datasets, we use YouTube Video Instance Segmentation [^2] https://youtube-vos.org/dataset/vis (YTVIS) the high quality version [^3] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which consists of daily-life diverse and complex videos downloaded from YouTube. This dataset is used for both object discovery and object recognition. We also include the synthetic video-text dataset CLEVRER [^4] http://clevrer.csail.mit.edu, which consists of geometric objects and paired question texts, testing the multi-modal reasoning capability. This dataset is used for visual question answering.
@@
Object discovery. We compare our method RandSF.Q with baselines STEVE, VideoSAUR and SlotContrast in terms of their object discovery performance on datasets MOVi-C, MOVi-D and YTVIS. This is basically an unsupervised object segmentation task. We use multiple recognized metrics for comprehensive measurement: Adjusted Rand Index (ARI) [^5] https://scikit-learn.org/stable/modules/generated/sklearn.metri
@@
As shown in Table 1, across all datasets and almost all metrics, our RandSF.Q defeats all the baselines. Specifically, on dataset YTVIS, RandSF.Q surpasses the latest SotA SlotContrast by more than 10 points by metrics ARI and ARIfg, and up to 6 points by mBO and mIoU. In contrast, SlotContrast only boost their baseline VideoSAUR limitedly, less than 3 points by metrics mBO and mIoU, or even degrading by ARI and ARIfg.
Object recognition. We compare our method RandSF.Q with baseline SlotContrast in terms of their object recognition performance on dataset YTVIS. This consists of two sub-tasks, i.e., supervised classification and regression from slots. Some literature (locatello2020slotattent) names this as set prediction. We follow (seitzer2023dinosaur) to represent the dataset as slots and train a two-layer MLP to predict the object class and bounding box corresponding to each slot, supervised by the annotations of object class labels and bounding boxes in the dataset. We use top1/top3 accuracy and box IoU score as the class label classification and bounding box regression performance metrics.
@@ Conclusion
Based on our two observations, we propose our method RandSF.Q. We realize informative query prediction by utilizing the next feature and effective query prediction learning by randomly sampling slot-feature pairs from available recurrences as transitioner inputs for query prediction. Our method show significant performance advantages over the latest SotA methods on scene representation and scene understanding. Our core claim is also verified through performance matrix analysis. Our work pushes the video OCL research greatly forward.
@@
Acknowledgments
We acknowledge the support of Finnish Center for Artificial Intelligence (FCAI), Research Council of Finland flagship program. We thank the Research Council of Finland for funding the projects ADEREHA (grant no. 353198), BERMUDA (362407) and PROFI7 (352788). We also appreciate CSC - IT Center for Science, Finland, for granting access to supercomputers Mahti and Puhti, as well as LUMI, owned by the European High Performance Computing Joint Undertaking (EuroHPC JU) and hosted by CSC Finland in collaboration with the LUMI consortium. Furthermore, we acknowledge the computational resources provided by the Aalto Science-IT project through the Triton cluster.